# On a conjecture of Peter Neumann on fixed points in permutation groups

### v1

@@ 1. Introduction
Most of the work in the literature on this conjecture has concerned the case of primitive groups of affine type. For these we have $G=V\rtimes G_{0}\leq\mathrm{AGL}(V)$, where $V$ is a finite vector space, and $G_{0}$ is an irreducible subgroup of $\mathrm{GL}(V)$; the conjecture says that such a group $G_{0}$ should possess an element $g$ such that $\dim C_{V}(g)\leq\frac{1}{2}\dim V$. For $G$ solvable, Neumann proved this, with the stronger upper bound $\frac{7}{18}\dim V$ instead of $\frac{1}{2}\dim V$; and he conjectured that the true bound in this case should be $\frac{1}{3}\dim V$, noting that equality is achieved by the group $G_{0}=A_{4}<\mathrm{GL}_{3}(p)$ for odd primes $p$. Of course Neumann did not have the CFSG available to use at the time. Using CFSG, Segal and Shalev [72, Lemma 2.3] proved the original Neumann conjecture for affine groups; following a further improvement in [34], the affine conjecture with the bound $\frac{1}{3}\dim V$ was finally established by Guralnick and Malle in [27]. Generalizing Neumann’s example, one can see that the bound is sharp for the groups $G_{0}=\mathrm{SO}_{3}(q)<\mathrm{GL}_{3}(q)$ for $q$ an odd prime power.
@@
If $G$ is not affine, we will in fact find $g\in G$ with $1\leq\mathrm{fix}(g)<n^{1/3}$; see Theorem 3.1, Proposition 3.6 and the proofs of Propositions 3.2 and 3.3.
@@ 2. Preliminaries
For $S=A_{n}$ with $n\geq 6$, we take $g$ to be either an $n$-cycle or an (n1)(n-1)-cycle; and for $S=A_{5}$ we take $g$ to be a $3$-cycle. For $S$ sporadic we use [12].
@@
then by considering at the order formula for $|G|$, we see that $|G|$ is divisible by a primitive prime divisor (ppd) of $q^{i}-1$ for some $i\geq 3$ with $i\neq 4$. Such a ppd is congruent to $1\pmod{i}$, and hence is at least $7$.
@@ 3. Reduction to simple groups
In this case we have $G\leq A\wr S_{t}$ acting on $\Omega=\Delta^{t}$, where $A$ is an almost simple or simple diagonal type primitive group on $\Delta$, and ${\rm Soc}(G)={\rm Soc}(A)^{t}$. Set $m=|\Delta|$. By the proof of Proposition 3.2 for the simple diagonal case, and by Theorem 3.1 for the almost simple case, there exists $x\in{\rm Soc}(A)$ such that $1\leq\mathrm{fix}(x,\Delta)<m^{1/3}$. Letting $g=(x,\ldots,x)\in{\rm Soc}(G)$, we then have $1\leq\mathrm{fix}(g,\Omega)<m^{t/3}=n^{1/3}$, as required. ∎
@@
Let $S$ be a finite non-abelian simple group. There exists an element $x\in S$ of prime order $p\geq 5$ such that for any $S$-set of size $n$ on which $S$ has no fixed points, the number of orbits of $\langle x\rangle$ is at most $n/3$. Moreover, we can choose $x\in S$ of prime power order such that the number of orbits of $\langle x\rangle$ is less than $n/3$ unless either $S=A_{5}$ and every orbit of has size $6$ or $12$, or $S=A_{6}$ and every orbit has size $6$.
@@
In case (i), $\langle x\rangle$ has at most $2n/(p+1)$ orbits, and since $p\geq 5$, this is at most $n/3$, giving the first assertion of the lemma. Moreover, we get the strict inequality as we long as we can choose $p>5$, and so by Lemma 2.4 we only need to consider $S=A_{5}$, $A_{6}$ or $U_{4}(2)$. In the last case, strict inequality holds for an element of order $5$ and in the first two cases, the strict inequality holds for an element of order $5$ apart from the case where $n=6$ and $S=A_{5}$ or $A_{6}$; for $S=A_{5}$, equality also holds for the transitive action of degree 12.
Consider case (iii). Inspection shows that in all but one case, we can choose a different element $x\in S$ of order $p$ that does not lie in [10, Table 6], hence satisfies (i). The exception is $S=\mathrm{PSL}_{2}(p+1)$ with $p$ a Mersenne prime. In this case, for all actions $(S,\Omega)$ we have $\mathrm{fix}(x,\Omega)\leq 2n/(p+2)$ (with equality for the action on the cosets of a Borel subgroup); for $p\geq 7$, this is at most $2/9$ and again, $\langle x\rangle$ has at most $n/3$ orbits. If $p>7$, strict inequality holds and if $p=7$ we choose an element of order $9$ and strict inequality holds (for all actions of $S$).
@@
    If $S=A_{5}$ or $A_{6}$, then there exists an element $g\in T$ of order $5$ such that one of the following holds:
@@
      each $T$-orbit has size $6$ or $12$ for $n=5$, or $6$ for $n=6$; in particular, the kernel of $T$ on each orbit contains all but one component of $T$.
@@
Suppose first that the projection of $H$ into at least two of the components of $T$ is non-surjective; say these projections are $M_{1}<S$ and $M_{2}<S$. We can replace $H$ by its overgroup $M_{1}\times M_{2}\times S^{k-2}$, so that the action is that of $S^{2}$ on the cosets of $M_{1}\times M_{2}$. For $i=1,2$ write $n_{i}=|S:M_{i}|$. Let $f_{i}$ be proportion of fixed points of $x$ on $S/M_{i}$. If $f_{1}f_{2}=0$ then $g$ is a derangement of odd prime power order at least $5$, so the number of orbits of $\langle g\rangle$ is less than $n/3$. Assume then $f_{1}f_{2}\neq 0$, so $f_{1}f_{2}<f_{1}$. If $|x|$ is prime, it follows at once that the number of orbits of $\langle g\rangle$ divided by $n$ is strictly less than the number of orbits of $\langle x\rangle$ on $S/M_{1}$ divided by $n_{1}$, and we are done by Lemma 3.4. Assume now $|x|$ is not prime; by inspecting the proof of Lemma 3.4, we have $|x|=9$ (and $S=A_{9},A_{10},\mathrm{PSL}_{2}(8)$) and the number of orbits of $\langle x\rangle$ on $S/M_{1}$ is strictly less than $n_{1}/3$. In particular, we see that the number of orbits of $\langle g\rangle$ is strictly less than $(n_{1}/3)n_{2}=n/3$, and we are done.
@@
Finally, suppose all projections are surjective. We show strict inequality in this case. Replacing $H$ by a maximal subgroup containing it, we are reduced to the case $k=2$ and $H\cong S$ is a diagonal subgroup. In this case, we have $\mathrm{fix}(g,\Omega)=|C_{S}(x)|$. We claim that $|C_{S}(x)|\leq|S|/12=n/12$. This is clear if $S$ has no subgroup of index at most $12$. Otherwise, $S$ embeds in $S_{12}$ and one can compute the list of such simple groups and check the claim directly. (Note that equality can occur for $x$ of order $5$ in $A_{5}$). Exclude the cases where $x$ is an element of order $9$ (i.e. $S=A_{9},A_{10}$ or $\mathrm{PSL}_{2}(8)$ – see the proof of Lemma 3.4). Thus $\mathrm{fix}(g,\Omega)\leq n/12$, and so the number of orbits of $\langle g\rangle$ is at most $n/12+11n/12p\leq(4/15)n$ and the result follows. In the three cases remaining, one computes directly that the result holds. ∎
@@
Let $A$ be a minimal normal subgroup of $H$ with $A=S^{e}$ for some nonabelian simple group $S$ and positive integer $e$. No orbit of $A$ on $\Delta$ is trivial and so by Lemma 3.5, there exists $g\in A$ (namely $g=(x,\ldots,x)$ with $x$ as in Lemma 3.4) such that the number of orbits of $\langle g\rangle$ on $\Delta$ is at most $t/3$. Then $|C_{E}(g)|\leq|E|^{1/3}$. Moreover, unless $S\cong A_{5}$ or $A_{6}$, the inequality is strict.
It remains to consider the cases where $S=A_{5}$ or $A_{6}$ and equality holds. Let $g$ be as above. Every $A$-orbit has size $6$, or possibly $12$ for $S=A_{5}$. We see that $\Delta$ is the disjoint union of the nontrivial orbits for each of the components of $A$. Since $A$ has no fixed points, it follows that $A$ is the unique minimal normal subgroup of $H$, and so $H$ acts faithfully on $A$ by conjugation.
Let $D$ be a component of $E$. In this paragraph we will show that either the statement of the proposition holds or $N_{A}(D)=C_{A}(D)$. By [3, Theorem C(1)], $N_{H}(D)/C_{H}(D)$ is isomorphic to a subgroup of $\mathrm{Aut}(D)$ containing all inner automorphisms. Since $N_{A}(D)/C_{A}(D)$ is a normal subgroup of $N_{H}(D)/C_{H}(D)$, it follows that if $N_{A}(D)\neq C_{A}(D)$ then $N_{A}(D)/C_{A}(D)$ contains a subgroup inducing the full group of inner automorphisms of $D$. Assume this is the case. By the choice of $g$, we have $|C_{E}(g)|=|E|^{1/3}$ and $\langle g\rangle$ has precisely $t/3$ orbits on the components of $E$. In particular, $g$ centralizes each component it normalizes. Let $S_{1}$ be the unique component of $A$ not normalizing $D$. Then $N_{A}(D)=M\times C$, where $M=N_{S_{1}}(D)$ and $C$ is the product of all components of $A$ different from $S_{1}$. Note that if $S=A_{6}$ then $M\cong A_{5}$, and if $S=A_{5}$ then $M\cong D_{10}$ or $C_{5}$. We claim that the action of $N_{A}(D)$ on $D$ factors through one component of $N_{A}(D)$. If this is the case, then a diagonal element $g$ of $A$ of order $5$ acts nontrivially on $D$, against our assumption. If $M\cong A_{5}$ then $N_{A}(D)$ is perfect, so $N_{A}(D)/C_{A}(D)\cong D$, which implies the claim (since a normal subgroup of $N_{A}(D)$ is a product of components). If $M\cong D_{10}$ or $C_{5}$, then the image of $M$ in $\mathrm{Aut}(D)$ is a solvable subgroup normalized by $D$. Such image must be trivial, that is, $M$ centralizes $D$. In particular, also in this case $N_{A}(D)/C_{A}(D)\cong D$ and the claim holds.
@@
Since the outer automorphism group of $S$ is a $2$-group, by viewing $w$ as an element of $\mathrm{Aut}(A)\cong\mathrm{Aut}(S)\wr S_{e}$ we see that there exists $a\in A$ such that $wa$ centralizes a diagonal element $g^{\prime}$ of $A$ of order $5$, and such that $wa$ is an $\ell$-element. Now replace $w$ by $wa$ and $g$ by $g^{\prime}$, and look at $z:=wg$, so $|C_{E}(z)|\leq|C_{E}(g)|\leq|E|^{1/3}$. It remains to prove that the inclusion $C_{E}(z)<C_{E}(g)$ is strict. The inclusion is strict if $z$ does not have the same orbits as $g$ on components of $E$, so assume that the orbits of $z$ and $g$ are the same. This implies that $w$ normalizes each component of $E$, and so $w$ does not centralize some component of $E$, say $L$. If $g$ normalizes $L$, then by the above argument $g$ centralizes $L$, and so $C_{E}(z)<C_{E}(g)$. If $g$ does not normalize $L$, the $g$-orbit of $L$ (in the action on $\Delta$) has size $5$, so $B:=\prod_{i}L^{g^{i}}\cong L^{5}$. We have that $C_{B}(g)$ is a diagonal subgroup isomorphic to $L$. Then $w$ does not centralize $C_{B}(g)$, so $C_{E}(z)<C_{E}(g)$ and this concludes the proof. ∎
@@ 4. Alternating groups
Assume $M=(S_{k}\wr S_{t})\cap G$, where $n=kt$ and $1<t<n$. If $n$ is odd, we choose $g$ an $n$-cycle, and then $\mathrm{fix}(g,G/M)=1$ (since $g$ must permute the blocks transitively and has a unique cycle). If $n$ is even and $k\neq 2$, we choose $g\in M$ with cycle type (1,1,k2,nk)(1,1,k-2,n-k), and we have $\mathrm{fix}(g,G/M)=1$: indeed, since $k-2\neq n-k$, the two fixed points must necessarily belong to the same block of any $k^{t}$-partition fixed by $g$, and then the (k2)(k-2)-cycle must preserve that block, and the other blocks are determined as $g$ acts as a cycle on their points. Finally, if $n$ is even and $k=2$, we choose $g\in M$ with cycle type (n2,2)(n-2,2). Since $t>2$, the $2$-cycle must necessarily stabilize a block, and the remaining blocks are determined since $g$ acts as a cycle on their points, so $\mathrm{fix}(g,G/M)=1$. ∎

### v2

@@ 1. Introduction
Most of the work in the literature on this conjecture has concerned the case of primitive groups of affine type. For these we have $G=V\rtimes G_{0}\leq\mathrm{AGL}(V)$, where $V$ is a finite vector space, and $G_{0}$ is an irreducible subgroup of $\mathrm{GL}(V)$; the conjecture says that such a group $G_{0}$ should possess an element $g$ such that $\dim C_{V}(g)\leq\frac{1}{2}\dim V$. For $G$ solvable, Neumann proved this, with the stronger upper bound $\frac{7}{18}\dim V$ instead of $\frac{1}{2}\dim V$; and he conjectured that the true bound in this case should be $\frac{1}{3}\dim V$, noting that equality is achieved by the group $G_{0}=A_{4}<\mathrm{GL}_{3}(p)$ for odd primes pp. Of course Neumann did not have the CFSG available to use at the time. Using CFSG, Segal and Shalev [72, Lemma 2.3] proved the original Neumann conjecture for affine groups; following a further improvement in [34], the affine conjecture with the bound $\frac{1}{3}\dim V$ was finally established by Guralnick and Malle in [27]. Generalizing Neumann’s example, one can see that the bound is sharp for the groups $G_{0}=\mathrm{SO}_{3}(q)<\mathrm{GL}_{3}(q)$ for $q$ an odd prime power.
@@
If $G$ is not affine, we will in fact find $g\in G$ with $1\leq\mathrm{fix}(g)<n^{1/3}$; see Theorem 3.1, Proposition 3.6 and the proofs of Propositions 3.2 and $3.$
@@ 2. Preliminaries
For $S=A_{n}$ with $n\geq 6$, we take $g$ to be either an $n$-cycle or an (n1)(n-1)-cycle; and for $S=A_{5}$ we take $g$ to be a 33-cycle. For $S$ sporadic we use [12].
@@
then by considering at the order formula for $|G|$, we see that $|G|$ is divisible by a primitive prime divisor (ppd) of $q^{i}-1$ for some $i\geq 3$ with $i\neq 4$. Such a ppd is congruent to $1\pmod{i}$, and hence is at least 77.
@@ 3. Reduction to simple groups
In this case we have $G\leq A\wr S_{t}$ acting on $\Omega=\Delta^{t}$, where AA is an almost simple or simple diagonal type primitive group on $\Delta$, and ${\rm Soc}(G)={\rm Soc}(A)^{t}$. Set $m=|\Delta|$. By the proof of Proposition 3.2 for the simple diagonal case, and by Theorem 3.1 for the almost simple case, there exists $x\in{\rm Soc}(A)$ such that $1\leq\mathrm{fix}(x,\Delta)<m^{1/3}$. Letting $g=(x,\ldots,x)\in{\rm Soc}(G)$, we then have $1\leq\mathrm{fix}(g,\Omega)<m^{t/3}=n^{1/3}$, as required. ∎
@@
Let $S$ be a finite non-abelian simple group. There exists an element $x\in S$ of prime order $p\geq 5$ such that for any $S$-set of size $n$ on which $S$ has no fixed points, the number of orbits of $\langle x\rangle$ is at most $n/3$. Moreover, we can choose $x\in S$ of prime power order such that the number of orbits of $\langle x\rangle$ is less than $n/3$ unless either $S=A_{5}$ and every orbit of has size 66 or $12$, or $S=A_{6}$ and every orbit has size 66.
@@
In case (i), $\langle x\rangle$ has at most $2n/(p+1)$ orbits, and since $p\geq 5$, this is at most $n/3$, giving the first assertion of the lemma. Moreover, we get the strict inequality as we long as we can choose $p>5$, and so by Lemma 2.4 we only need to consider $S=A_{5}$, $A_{6}$ or $U_{4}(2)$. In the last case, strict inequality holds for an element of order 55 and in the first two cases, the strict inequality holds for an element of order 55 apart from the case where $n=6$ and $S=A_{5}$ or $A_{6}$; for $S=A_{5}$, equality also holds for the transitive action of degree 12.
Consider case (iii). Inspection shows that in all but one case, we can choose a different element $x\in S$ of order $p$ that does not lie in [10, Table 6], hence satisfies (i). The exception is $S=\mathrm{PSL}_{2}(p+1)$ with $p$ a Mersenne prime. In this case, for all actions $(S,\Omega)$ we have $\mathrm{fix}(x,\Omega)\leq 2n/(p+2)$ (with equality for the action on the cosets of a Borel subgroup); for $p\geq 7$, this is at most $2/9$ and again, $\langle x\rangle$ has at most $n/3$ orbits. If $p>7$, strict inequality holds and if $p=7$ we choose an element of order 99 and strict inequality holds (for all actions of $S$).
@@
    If $S=A_{5}$ or $A_{6}$, then there exists an element $g\in T$ of order 55 such that one of the following holds:
@@
      each $T$-orbit has size 66 or $12$ for $n=5$, or 66 for $n=6$; in particular, the kernel of $T$ on each orbit contains all but one component of $T$.
@@
Suppose first that the projection of $H$ into at least two of the components of $T$ is non-surjective; say these projections are $M_{1}<S$ and $M_{2}<S$. We can replace $H$ by its overgroup $M_{1}\times M_{2}\times S^{k-2}$, so that the action is that of $S^{2}$ on the cosets of $M_{1}\times M_{2}$. For $i=1,2$ write $n_{i}=|S:M_{i}|$. Let $f_{i}$ be proportion of fixed points of $x$ on $S/M_{i}$. If $f_{1}f_{2}=0$ then $g$ is a derangement of odd prime power order at least 55, so the number of orbits of $\langle g\rangle$ is less than $n/3$. Assume then $f_{1}f_{2}\neq 0$, so $f_{1}f_{2}<f_{1}$. If $|x|$ is prime, it follows at once that the number of orbits of $\langle g\rangle$ divided by $n$ is strictly less than the number of orbits of $\langle x\rangle$ on $S/M_{1}$ divided by $n_{1}$, and we are done by Lemma 3.4. Assume now $|x|$ is not prime; by inspecting the proof of Lemma 3.4, we have $|x|=9$ (and $S=A_{9},A_{10},\mathrm{PSL}_{2}(8)$) and the number of orbits of $\langle x\rangle$ on $S/M_{1}$ is strictly less than $n_{1}/3$. In particular, we see that the number of orbits of $\langle g\rangle$ is strictly less than $(n_{1}/3)n_{2}=n/3$, and we are done.
@@
Finally, suppose all projections are surjective. We show strict inequality in this case. Replacing $H$ by a maximal subgroup containing it, we are reduced to the case $k=2$ and $H\cong S$ is a diagonal subgroup. In this case, we have $\mathrm{fix}(g,\Omega)=|C_{S}(x)|$. We claim that $|C_{S}(x)|\leq|S|/12=n/12$. This is clear if $S$ has no subgroup of index at most $12$. Otherwise, $S$ embeds in $S_{12}$ and one can compute the list of such simple groups and check the claim directly. (Note that equality can occur for $x$ of order 55 in $A_{5}$). Exclude the cases where $x$ is an element of order 99 (i.e. $S=A_{9},A_{10}$ or $\mathrm{PSL}_{2}(8)$ – see the proof of Lemma 3.4). Thus $\mathrm{fix}(g,\Omega)\leq n/12$, and so the number of orbits of $\langle g\rangle$ is at most $n/12+11n/12p\leq(4/15)n$ and the result follows. In the three cases remaining, one computes directly that the result holds. ∎
@@
Let AA be a minimal normal subgroup of $H$ with $A=S^{e}$ for some nonabelian simple group $S$ and positive integer $e$. No orbit of AA on $\Delta$ is trivial and so by Lemma 3.5, there exists $g\in A$ (namely $g=(x,\ldots,x)$ with $x$ as in Lemma 3.4) such that the number of orbits of $\langle g\rangle$ on $\Delta$ is at most $t/3$. Then $|C_{E}(g)|\leq|E|^{1/3}$. Moreover, unless $S\cong A_{5}$ or $A_{6}$, the inequality is strict.
It remains to consider the cases where $S=A_{5}$ or $A_{6}$ and equality holds. Let $g$ be as above. Every AA-orbit has size 66, or possibly $12$ for $S=A_{5}$. We see that $\Delta$ is the disjoint union of the nontrivial orbits for each of the components of AA. Since AA has no fixed points, it follows that AA is the unique minimal normal subgroup of $H$, and so $H$ acts faithfully on AA by conjugation.
Let $D$ be a component of $E$. In this paragraph we will show that either the statement of the proposition holds or $N_{A}(D)=C_{A}(D)$. By [3, Theorem C(1)], $N_{H}(D)/C_{H}(D)$ is isomorphic to a subgroup of $\mathrm{Aut}(D)$ containing all inner automorphisms. Since $N_{A}(D)/C_{A}(D)$ is a normal subgroup of $N_{H}(D)/C_{H}(D)$, it follows that if $N_{A}(D)\neq C_{A}(D)$ then $N_{A}(D)/C_{A}(D)$ contains a subgroup inducing the full group of inner automorphisms of $D$. Assume this is the case. By the choice of $g$, we have $|C_{E}(g)|=|E|^{1/3}$ and $\langle g\rangle$ has precisely $t/3$ orbits on the components of $E$. In particular, $g$ centralizes each component it normalizes. Let $S_{1}$ be the unique component of AA not normalizing $D$. Then $N_{A}(D)=M\times C$, where $M=N_{S_{1}}(D)$ and $C$ is the product of all components of AA different from $S_{1}$. Note that if $S=A_{6}$ then $M\cong A_{5}$, and if $S=A_{5}$ then $M\cong D_{10}$ or $C_{5}$. We claim that the action of $N_{A}(D)$ on $D$ factors through one component of $N_{A}(D)$. If this is the case, then a diagonal element $g$ of AA of order 55 acts nontrivially on $D$, against our assumption. If $M\cong A_{5}$ then $N_{A}(D)$ is perfect, so $N_{A}(D)/C_{A}(D)\cong D$, which implies the claim (since a normal subgroup of $N_{A}(D)$ is a product of components). If $M\cong D_{10}$ or $C_{5}$, then the image of $M$ in $\mathrm{Aut}(D)$ is a solvable subgroup normalized by $D$. Such image must be trivial, that is, $M$ centralizes $D$. In particular, also in this case $N_{A}(D)/C_{A}(D)\cong D$ and the claim holds.
@@
Since the outer automorphism group of $S$ is a 22-group, by viewing $w$ as an element of $\mathrm{Aut}(A)\cong\mathrm{Aut}(S)\wr S_{e}$ we see that there exists $a\in A$ such that wawa centralizes a diagonal element $g^{\prime}$ of AA of order 55, and such that wawa is an $\ell$-element. Now replace $w$ by wawa and $g$ by $g^{\prime}$, and look at $z:=wg$, so $|C_{E}(z)|\leq|C_{E}(g)|\leq|E|^{1/3}$. It remains to prove that the inclusion $C_{E}(z)<C_{E}(g)$ is strict. The inclusion is strict if $z$ does not have the same orbits as $g$ on components of $E$, so assume that the orbits of $z$ and $g$ are the same. This implies that $w$ normalizes each component of $E$, and so $w$ does not centralize some component of $E$, say $L$. If $g$ normalizes $L$, then by the above argument $g$ centralizes $L$, and so $C_{E}(z)<C_{E}(g)$. If $g$ does not normalize $L$, the $g$-orbit of $L$ (in the action on $\Delta$) has size 55, so $B:=\prod_{i}L^{g^{i}}\cong L^{5}$. We have that $C_{B}(g)$ is a diagonal subgroup isomorphic to $L$. Then $w$ does not centralize $C_{B}(g)$, so $C_{E}(z)<C_{E}(g)$ and this concludes the proof. ∎
@@ 4. Alternating groups
Assume $M=(S_{k}\wr S_{t})\cap G$, where $n=kt$ and $1<t<n$. If $n$ is odd, we choose $g$ an $n$-cycle, and then $\mathrm{fix}(g,G/M)=1$ (since $g$ must permute the blocks transitively and has a unique cycle). If $n$ is even and $k\neq 2$, we choose $g\in M$ with cycle type (1,1,k2,nk)(1,1,k-2,n-k), and we have $\mathrm{fix}(g,G/M)=1$: indeed, since $k-2\neq n-k$, the two fixed points must necessarily belong to the same block of any $k^{t}$-partition fixed by $g$, and then the (k2)(k-2)-cycle must preserve that block, and the other blocks are determined as $g$ acts as a cycle on their points. Finally, if $n$ is even and $k=2$, we choose $g\in M$ with cycle type (n2,2)(n-2,2). Since $t>2$, the 22-cycle must necessarily stabilize a block, and the remaining blocks are determined since $g$ acts as a cycle on their points, so $\mathrm{fix}(g,G/M)=1$. ∎